A survey by the Canadian Frequent Flyers association revealed that 60% of its members made independent airline reservations last year. If two members are selected at random, what is the probability that,
a)Both made reservations last year?
b)at least one made reservations last year?
c)only one made a reservation last year?

Answer by stanbon(75887)

60% of its members made independent airline reservations last year. If two members are selected at random, what is the probability that,
a)Both made reservations last year?
Prob = 0.6*0.6 = 0.36
-----------------
b)at least one made reservations last year?
P(at least one) = 1- P(none made a reservation)
= 1- 0.4*0.4
= 1 - 0.16
=0.84
-------------
c)only one made a reservation last year?
Prob = 2C1 (0.6)(0.4)
=2*0.24
=0.48
---------
Cheers,
Stan H.